RC meeting and Run Plan, Tuesday Oct. 16
- Beam Quality Bottom line:
- Beam looked great after beam tuning operations up to 55uA
- Ops still committed to get us the remaining 15uA to go to 70uA
- Beam was steady for luminosity scan
- Hopefully things will keep looking good after RF recovery
- Finished luminosity scan with flying colors!
Activities during opportunistic controlled access
- Brad and Pavel:
- Installed NDX detector in SHMS hut (<1hour)
- Steve Lassiter and Andy Kenyon:
- Lock Hall back up ASAP to show we are ready for beam.:
- Called MCC at noon
- SHMS at small angle
- Almost done running at 6.49 degrees, didn't encounter issues with the diffusor ion chambers
- We understand rates enough to avoid DC trips
- During Luminosity scan found that DC trips start happening around 800-900kHz singles rate
- At this point, we are either just below, or just above Itrip values for the DC cathodes.
- Took some data with slightly higher Itrip value (endorsed by Rolf and Mark) during luminosity scan to better understand SHMS performance near the high-rate limit.
- Reached rates up to ~1MHz
Issues: Requiring attention
- Still bringing down HMS Dipole LN2 level to below yellow alarm level
- Switched to manual mode because auto-mode was happy to stay at 90+%
- Maybe increase upper yellow and red alarm levels if we aren't worried about them?
- Hall B asked if they could remove halfway plate until Monday
- We agreed.
- Hall D needs opportunistic 1shift of HallC down for beam studies, not pressing right now (needs to be done before end-of-run)
- Might be possible when swapping out aerogel trays for the next experiment.
- Greg Smith: Dave Meekins needs access to work on deuterium target for next experiment
- Called today after I locked Hall back up
- Told him there will be opportunity when swapping out aerogel on Monday
Plans - Physics
- Run 2-3 more shifts with Q2=5.5. W=3.02 x=0.40 theta_Shms=6.49
- Currently collected 26% of statistics + Al data in 3.25 shifts
- Can run up to 45uA in this setting
- Wednesday-Friday: Run 5 shifts with Q2=5.5. W=3.02 x=0.40 theta_Shms=9.56
- Collect at least 50% of statistics goal
- Friday-Saturday: Run 4-5 shifts with Q2=5.5. W=3.02 x=0.40 theta_Shms=12.56
- Match statistics for the theta_Shms=6.49 point
- FINALLY: 4-6 shift buffer
- Is this enough buffer?
- Wishlist of settings to revisit?